A flat board leans against a wall. The upper edge is 6 feet above the floor and the lower edge is 2 feet out from the wall. What is the slope of the board?
Answer by ABU1998(5)

slope=rise(the y-value)/run (the x-value)
if the flat board leans from left to right (the wall is on the right) then the slope is positive,slope = 6/2=3.if the wall on the left then the slope is negative slope = -6/2=-3.hope it helps